# Selective Mixup for Debiasing Question Selection in Computerized Adaptive Testing

### v1

@@ Abstract.
Computerized Adaptive Testing is a widely used technology for evaluating learners’ proficiency in online education platforms. By leveraging prior estimates of proficiency to select questions and updating the estimates iteratively based on responses, CAT enables personalized learner modeling and has attracted substantial attention. Despite this progress, most existing works focus primarily on improving diagnostic accuracy, while overlooking the selection bias inherent in the adaptive process. Selection Bias arises because the question selection is strongly influenced by the estimated proficiency, such as assigning easier questions to learners with lower proficiency and harder ones to learners with higher proficiency. Since the selection depends on prior estimation, this bias propagates into the diagnosis model, which is further amplified during iterative updates, leading to misalignment and biased predictions. Moreover, the imbalanced nature of learners’ historical interactions often exacerbates the bias in diagnosis models. To address this issue, we propose a debiasing framework consisting of two key modules: Cross-Attribute Examinee Retrieval and Selective Mixup-based Regularization. First, we retrieve balanced examinees with relatively even distributions of correct and incorrect responses and use them as neutral references for biased examinees. Then, mixup is applied between each biased examinee and its matched balanced counterpart under label consistency. This augmentation enriches the diversity of bias-conflicting samples and smooths selection boundaries. Finally, extensive experiments on two benchmark datasets with multiple advanced diagnosis models demonstrate that our method substantially improves both the generalization ability and fairness of question selection in CAT.
Computerized Adaptive Testing, Intelligent Education, Bias Mitigation
@@ 1. Introduction
Figure 1. Training Process of Data-Driven Question Selection Module.
Computerized Adaptive Testing (CAT) enables accurate proficiency estimation with fewer questions by iteratively selecting questions and updating learner states (liu2024survey). While early CAT relied on heuristic strategies such as Maximum Fisher Information (lord2012mfi), recent advances have shifted toward data-driven approaches that learn the Question Selection Module from large-scale interaction data (ghosh2021bobcat; wang2023gmocat; yu2024ucats), achieving strong diagnostic performance. However, research on mitigating bias in CAT remains limited (kwon2023useraif; hort2024biassurvey). Bias arising from data imbalance or adaptive selection may lead to systematically easier or harder questions being assigned, thereby distorting proficiency estimates and undermining fairness. Prior work such as UserAIF (kwon2023useraif) alleviates bias by recalibrating question parameters, but it overlooks the effect of biased training data on the Question Selection Module itself.
The origins and consequences of biased selection remain insufficiently explored, leaving a gap in the development of fair and robust CAT systems. This work addresses the gap by examining how biased selection arises during training and affects both accuracy and fairness, with a focus on mitigating bias within data-driven Question Selection Modules. Recent Question Selection Modules often adopt a bi-level optimization framework (ghosh2021bobcat; yu2024ucats), where each examinee’s interactions are split into a support set and a meta set. The support set is used by the Cognitive Diagnosis Module to estimate proficiency, while the meta set provides feedback for training the Question Selection Module. As illustrated in Figure 1, at each step $t$, the Question Selection Module selects the next question based on the examinee’s history, the CDM updates proficiency after the response, and predictions on the meta set yield feedback to optimize the Question Selection Module. This process repeats for $T$ steps to iteratively refine question selection.
In this process, the meta set guides how the Question Selection Module is trained, aiming to align the distribution of selected questions with that of the meta set to improve diagnostic accuracy. In practice, however, the meta set is inherently imbalanced. As training progresses, the Question Selection Module overfits these skewed patterns, causing a divergence between selected and unbiased distributions, which exacerbates selection bias.
@@
Figure 2. The imbalanced distribution in the meta set can lead to a shift in the selection set, resulting in biased estimates of examinee proficiency.
Examinees differ in their proportion of correct responses: some perform well, others poorly, while some are more balanced. We characterize each examinee by this correct-response ratio. As shown in Figure 2(a), the distribution of meta sets and selected data in the NIPS-EDU dataset varies with this ratio [^1] https://eedi.com/projects/neurips-education-challenge, with density curves estimated by histograms and cubic spline smoothing. For high-ratio examinees, the meta set centers around 0.6 but shifts toward 1.0 after training, indicating a bias toward reinforcing strong performance. For low-ratio examinees, the distribution similarly shifts toward 0.0, amplifying underperformance. Thus, an already imbalanced meta set becomes more skewed during training. This imbalance introduces selection bias: the Question Selection Module tends to favor easier questions for high-ratio examinees and harder ones for low-ratio examinees, regardless of their true proficiency. While data-driven CAT models achieve strong empirical performance, such biased behavior contradicts the core objective of CAT and leads to distorted outcomes. As illustrated in Figure 2(b), this results in unreliable estimates, systematically overestimating high performers and underestimating low performers.
Selection bias in data-driven CAT arises from imbalanced meta sets, where rare cases—such as correct responses from low-performing students or incorrect responses from high-performing students—are underrepresented. This imbalance leads the Question Selection Module to overfit common patterns and hinders generalization. Since CAT jointly optimizes the Cognitive Diagnosis Module and the Question Selection Module, it is difficult to isolate the source of bias. To focus on selection bias, we fix the Cognitive Diagnosis Module and address imbalance in the Question Selection Module by augmenting rare cases with interactions from balanced examinees with matching labels. Through selective mixup of these samples, we construct a more balanced training distribution that guides the Question Selection Module to learn smoother, less biased decision boundaries, thereby improving robustness and fairness under unbiased evaluation.
  • Formally identify and analyze the problem of selection bias in data-driven CAT, tracing its origin to imbalanced meta set distributions during training.
  • Propose a selective cross-attribute mixup strategy to augment underrepresented interactions and enhance the generalization of the Question Selection Module.
  • Validate the effectiveness of the proposed method through extensive experiments on two benchmark datasets, demonstrating strong bias-mitigation capability .
@@ 2.1. CAT and Bias in Cognitive Modeling
Computerized Adaptive Testing (CAT) adaptively selects questions to estimate examinees’ knowledge mastery (chang2015survey; liu2024survey), typically consisting of a Cognitive Diagnosis Module (CDM) to infer knowledge states and a Question Selection Module to determine the next qustion. Early CDMs were built on Item Response Theory  (embretson2013irt) and its multidimensional extension  (ackerman2014mirt), while recent deep learning models such as NCDM (wang2020ncd), KaNCD (wang2022kancd), RCD (gao2021rcd), and DCD (chen2023dcd) have greatly improved representation and prediction. For Question Selection Module, heuristic approaches such as Maximum Fisher Information (lord2012mfi), KL-based selection  (chang1996kli), MAAT (bi2020maat), and BECAT (zhuang2023becat) are simple and interpretable but limited in modeling complex interactions. More recent data-driven methods leverage bi-level optimization or reinforcement learning, including BOBCAT (ghosh2021bobcat), NCAT (zhuang2022ncat), GMOCAT (wang2023gmocat), LACAT (cheng2024lacat), and UATS (yu2024ucats). However, data-driven methods often overlook bias in the data and selection process, which may compromise robustness in CAT.
@@
Bias has been increasingly recognized in CDMs, where predictions may differ systematically across sensitive groups such as gender or socioeconomic status (zhang2024psrcf; zhang2024faircd; xu2025fairwisa). Methods like adversarial learning or causal reasoning have been proposed to mitigate such disparities. However, bias mitigation in CAT remains relatively underexplored, and research on bias mitigation in CAT is still very limited. In practice, CAT interaction data is inherently subject to selection bias. Training data-driven CAT models directly on such biased data can distort parameter estimation and compromise the validity of the system. Prior work such as UserAIF (kwon2023useraif) uses influence functions to identify less biased samples and recalibrate the question parameters in Cognitive Diagnosis Module.
@@
Our distinction. Previous studiy focuses on question-side bias, while our work targets the Question Selection Module, mitigating bias from adaptive selection to improve fairness and robustness in CAT, directly correcting bias at its source rather than applying debiasing after question selection.
2.2. Bias Mitigation Methods
@@
While Machine Learning systems can relieve humans from tedious tasks and perform complex calculations at high speed, they are only as reliable as the data on which they are trained (barocas2016big). Since algorithms are not intentionally designed to incorporate bias, they often replicate or even amplify the spurious correlations present in real-world data, raising concerns about fairness and accountability (pedreshi2008discrimination; van2022overcoming). To address these issues, a variety of strategies have been proposed. Invariant Risk Minimization (IRM) seeks environment-invariant predictors (arjovsky2019irm), while GroupDRO minimizes worst-group risk but requires group labels (sagawa2019groupDRO). Reweighting strategies adjust sample importance to counter imbalance (japkowicz2000reweight), yet they only reshape optimization weights without enriching data diversity. In contrast, Mixup-based methods (zhang2018mixup) regularize models by interpolating samples and labels, with extensions such as Manifold Mixup (verma2019manifold) and Co-Mixup (kim2021co) improving hidden representation learning and pair selection. Building on this line of work, our method adopts a selective mixup strategy to mitigate selection bias in CAT, leveraging the simplicity and effectiveness of Mixup to enhance generalization.
@@ 3.1. Computerized Adaptive Testing
CAT reduces the burden on examinees by minimizing the number of questions while ensuring that the estimated proficiency $\boldsymbol{\theta}$ aligns with the true proficiency $\boldsymbol{\theta}_{0}$. It consists of a Cognitive Diagnosis Module, which estimates proficiency from responses, and a Question Selection Module, which adaptively selects the next item. Their feedback loop enables efficient and accurate assessment.
@@ 3.1.1. Cognitive Diagnosis Module
The Cognitive Diagnosis Module (CDM) estimates an examinee’s proficiency $\boldsymbol{\theta}$ by fitting their historical responses and predicting performance on unseen questions. Formally, the model outputs the probability $p\in[0,1]$ that an examinee answers a given item correctly. A representative neural network–based CDM is NCDM (wang2020ncd), which leverages neural networks to capture complex interactions between examinees and items. The model incorporates several components: an examinee proficiency vector $\boldsymbol{\theta}\in(0,1)^{1\times K}$, a question concept vector $\boldsymbol{Q}_{i}\in{0,1}^{1\times K}$ indicating the required knowledge concepts, a knowledge difficulty vector $\boldsymbol{h}^{diff}\in(0,1)^{1\times K}$, a scalar discrimination parameter $h^{disc}\in(0,1)$, and other trainable neural network parameters.
@@
The first interaction layer is inspired by the Multidimensional question Response Theory (MIRT) (ackerman2014mirt) model, and is defined as:
@@
where $\circ$ denotes the element-wise product.The intermediate layers consist of two fully connected layers followed by an output layer:
(2) $\boldsymbol{f}_{1}$ $=\varphi(\boldsymbol{W}_{1}\boldsymbol{x}^{T}+\boldsymbol{b}_{1}),$
(3) $\boldsymbol{f}_{2}$ $=\varphi(\boldsymbol{W}_{2}\boldsymbol{f}_{1}+\boldsymbol{b}_{2}),$
(4) $p$ $=\varphi(\boldsymbol{W}_{3}\boldsymbol{f}_{2}+\boldsymbol{b}_{3}),$
where $\varphi$ denotes the activation function, typically implemented as a sigmoid. The monotonicity assumption requires that the probability of correctly answering an item increases monotonically with respect to any dimension of a student’s knowledge proficiency. To enforce this property, NCDM adopts a simple yet effective strategy by constraining all elements of $\boldsymbol{W}_{1}$, $\boldsymbol{W}_{2}$, and $\boldsymbol{W}_{3}$ to be positive.
@@
The cross-entropy loss is adopted between the predicted probability $M(q;\boldsymbol{\theta},\zeta_{q},\zeta_{n})=p\in[0,1]$ and the ground-truth response $y$:
(5) $\ell(y,M(q;\boldsymbol{\theta},\zeta_{q},\zeta_{n}))=-\big(y\log p+(1-y)\log(1-p)\big),$
where $q$ denotes the exercise, $\boldsymbol{\theta}$ represents the examinee proficiency, $\zeta_{q}$ corresponds to the question parameters (e.g., concept vector, difficulty, discrimination), and $\zeta_{n}$ denotes the trainable neural network parameters. After training, the estimated proficiency vector $\boldsymbol{\theta}$ serves as the diagnostic result, representing the examinee’s proficiency over the $K$ knowledge concepts.
@@ 3.1.2. Question Selection Module
Existing data-driven Question Selection Module typically employ a bi-level optimization framework (ghosh2021bobcat), where a selection network is trained on large-scale examinee data. The dataset $\mathcal{D}$ is divided into a support set $\mathcal{D}_{u}$ and a meta set $\mathcal{D}_{m}$ used to compute the loss or reward. To illustrate, we briefly introduce BOBCAT (ghosh2021bobcat) as an example. At each step $t$, the selection network $\pi(\cdot)$ chooses a question $q_{t}\in\mathcal{D}_{u}$ based on the examinee’s current state $\boldsymbol{s_{t}}$, which is encoded from their interaction history $r=\{(q_{1},y_{1}),\ldots,(q_{t-1},y_{t-1})\}$, where $y_{i}\in\{0,1\}$ indicates whether the response was correct or incorrect.
The state $\boldsymbol{s_{t}}$ is a vector of length equal to the number of questions, where each entry is in $\{-1,0,1\}$, representing incorrect, unattempted, and correct responses, respectively. Formally, the selection process is defined as:
(6) $q_{t}=\pi(\boldsymbol{s_{t}},\mathcal{D}_{u};\zeta_{s})$
@@
To train the selection network, a bi-level optimization framework is adopted. The outer loop focuses exclusively on updating the selection network $\zeta_{s}$, since the Question Selection Module is assumed to be the main source of bias. The Cognitive Diagnosis Module is kept fixed, with all parameters $(\zeta_{q},\zeta_{n})$ pre-trained and frozen, while only the proficiency representation $\boldsymbol{\theta}$ is updated during inner optimization. After training, the selection network generates the next question for each examinee based solely on their current response state. The overall bi-level optimization procedure for data-driven question selection is summarized as follows:
@@
Given a selected question sequence, the examinee’s proficiency $\boldsymbol{\theta}_{i}$ is estimated by minimizing the prediction loss on their responses in $\mathcal{D}_{u}$ with fixed question parameters $(\zeta_{q},\zeta_{n})$, as defined in Eq. (5). Through this optimization, $\boldsymbol{\theta}_{i}$ is updated to obtain the estimated proficiency $\boldsymbol{\theta}_{i}^{*}$:
@@
where $T$ denotes the time step, i.e., the number of selected questions.
@@
The selection network $\zeta_{s}$ is then updated by minimizing the loss on the meta set $\mathcal{D}_{m}$ with the fixed proficiency $\boldsymbol{\theta}_{i}^{*}$ and pre-trained CDM parameters $(\zeta_{q},\zeta_{n})$:
@@
where $N$ denotes the number of examinees.

### v2

@@ Abstract.
Computerized Adaptive Testing is a widely used technology for evaluating examinees’ proficiency in online education platforms. By leveraging prior estimates of proficiency to select questions and updating the estimates iteratively based on responses, it enables personalized examinee modeling and has attracted substantial attention. Despite this progress, most existing works focus primarily on improving proficiency estimation accuracy, while overlooking the selection bias inherent in the adaptive process. Selection bias arises because the question selection is strongly influenced by the estimated proficiency, such as assigning easier questions to examinees with lower proficiency and harder ones to examinees with higher proficiency. Since the selection depends on prior estimation, this bias propagates into the diagnostic model, which is further amplified during iterative updates, leading to misaligned and biased predictions. Moreover, the imbalance in examinees’ historical interactions often exacerbates bias in diagnostic models. To address this issue, we propose a debiasing framework consisting of two key modules: Cross-Attribute Examinee Retrieval and Selective Mixup-based Regularization. First, we retrieve balanced examinees with relatively even distributions of correct and incorrect responses and use them as neutral references for biased examinees. Then, Mixup is applied between each biased examinee and its matched balanced counterpart under label consistency. This augmentation enriches the diversity of bias-conflicting samples and smooths selection boundaries. Finally, extensive experiments on two benchmark datasets with multiple advanced diagnosis models have been conducted. The results demonstrate that our method substantially improves the generalization ability of question selection.
Computerized Adaptive Testing, Bias Mitigation
@@ 1. Introduction
Figure 1. The training process of the data-driven Question Selection Module involves iterative updates guided by proficiency estimation and meta set feedback.
Computerized Adaptive Testing (CAT) enables accurate proficiency estimation with fewer questions by iteratively selecting questions and updating examinee states (liu2024survey; choi2020development; yang2022adaptive). While early CAT relied on heuristic strategies such as Maximum Fisher Information (lord2012mfi), recent advances have shifted toward data-driven approaches that learn the Question Selection Module from large-scale interaction data (ghosh2021bobcat; wang2023gmocat; yu2024ucats), achieving strong diagnostic performance. However, research on mitigating bias in CAT remains limited (kwon2023useraif; hort2024biassurvey). Bias arising from data imbalance or adaptive selection may lead to the systematic assignment of easier or harder questions, thereby distorting proficiency estimates and compromising fairness. Prior work such as UserAIF (kwon2023useraif) alleviates bias by recalibrating question parameters, but it overlooks the effect of biased training data on the Question Selection Module itself.
The origins and consequences of biased selection remain insufficiently explored, leaving a gap in the development of fair and robust CAT systems. This work addresses this limitation by examining how biased selection arises during training and impacts both proficiency estimation accuracy and fairness, with a focus on mitigating bias within data-driven Question Selection Modules. Recent Question Selection Modules often adopt a bi-level optimization framework (ghosh2021bobcat; yu2024ucats), where each examinee’s interactions are split into a support set and a meta set. The support set is used by the Cognitive Diagnosis Module (CDM) to estimate proficiency, while the meta set provides feedback for training the Question Selection Module. As illustrated in Figure 1, at each step $t$, the Question Selection Module selects the next question based on the examinee’s history, the CDM updates proficiency after the response, and predictions on the meta set yield feedback to optimize the Question Selection Module. This process repeats for $T$ steps to iteratively refine question selection.
In this process, the meta set guides how the Question Selection Module is trained, aiming to align the distribution of selected questions with that of the meta set to improve proficiency estimation accuracy. In practice, however, the meta set is inherently imbalanced. As training progresses, the Question Selection Module overfits these skewed patterns, leading to divergence between the selected and the unbiased distributions, thereby exacerbating selection bias.
@@
Figure 2. The imbalanced distribution of the meta set can induce shifts in the selection set, leading to biased estimates of examinee proficiency.
Examinees differ in the proportion of their correct responses: some perform well, others poorly, while some are more balanced. We characterize each examinee by this correct-response ratio. As shown in Figure 2(a), the distribution of meta sets and selected data in NIPS-EDU dataset varies with this ratio [^1] https://eedi.com/projects/neurips-education-challenge, with density curves estimated by histograms and cubic spline smoothing. For high-ratio examinees, the meta set centers around 0.6 but shifts toward 1.0 after training, indicating a bias toward reinforcing strong performance. For low-ratio examinees, the distribution similarly shifts toward 0.0, reinforcing underperformance. Thus, an already imbalanced meta set becomes more skewed during training. This imbalance introduces selection bias: the Question Selection Module tends to favor easier questions for high-ratio examinees and harder ones for low-ratio examinees, regardless of their true proficiency. While data-driven CAT models achieve strong empirical performance, such biased behavior contradicts the core objective of CAT and leads to distorted outcomes. As illustrated in Figure 2(b), this results in unreliable estimates, systematically overestimating high performers while underestimating low performers.
Selection bias in data-driven CAT arises from imbalanced meta sets, where rare cases such as correct responses from low-ratio examinees or incorrect responses from high-ratio examinees are underrepresented. This imbalance leads the Question Selection Module to overfit common patterns and weakens generalization. Since CAT jointly optimizes the CDM and the Question Selection Module, isolating the source of bias is challenging. To focus on selection bias, we fix the CDM and mitigate imbalance in the Question Selection Module by augmenting rare cases with interactions from balanced examinees sharing the same labels. Through selective Mixup of these samples, we obtain a more balanced and representative training distribution that guides the Question Selection Module toward smoother and less biased decision boundaries. Our contributions are summarized as follows:
  • Formally identify and analyze the problem of selection bias in data-driven CAT, tracing its origin to imbalanced meta-set distributions during training.
  • Propose a selective cross-attribute Mixup strategy that augments underrepresented interactions, enhances training diversity, and strengthens the generalization of the Question Selection Module.
  • Validate the proposed method through extensive experiments on two benchmark datasets, demonstrating its strong capability in mitigating bias.
@@ 2.1. CAT and Bias in Cognitive Modeling
Computerized Adaptive Testing (CAT) adaptively selects questions to estimate examinees’ knowledge proficiency (chang2015survey; liu2024survey), typically consisting of a Cognitive Diagnosis Module (CDM) to infer proficiency and a Question Selection Module to determine the next question. Early CDMs were built on Item Response Theory (embretson2013irt) and its multidimensional extension (ackerman2014mirt), while recent deep learning models such as NCDM (wang2020ncd), KaNCD (wang2022kancd), RCD (gao2021rcd), and DCD (chen2023dcd) have greatly improved representation and proficiency estimation accuracy. For the Question Selection Module, heuristic approaches such as Maximum Fisher Information (lord2012mfi), KL-based selection (chang1996kli), MAAT (bi2020maat), RAT (zhuang2022robust), and BECAT (zhuang2023becat) are simple and interpretable, but limited in capturing complex interactions. More recent data-driven methods leverage bi-level optimization or reinforcement learning, including BOBCAT (ghosh2021bobcat), NCAT (zhuang2022ncat), GMOCAT (wang2023gmocat), LACAT (cheng2024lacat), and UATS (yu2024ucats). However, data-driven methods often overlook bias in the data and selection process, which may compromise robustness in CAT.
@@
Bias has been increasingly recognized in Cognitive Modeling, where predictions may differ systematically across sensitive attributes such as gender or socioeconomic status (xu2025fairwisa). Methods like adversarial learning or causal reasoning (zhang2024faircd; zhang2024psrcf) have been proposed to mitigate such disparities. However, bias mitigation in CAT remains underexplored. In practice, CAT interaction data is inherently subject to selection bias, and directly training data-driven CAT models on such biased data can distort proficiency estimation and compromise system validity. Prior work such as UserAIF (kwon2023useraif) uses influence functions to identify less biased samples and recalibrate the question parameters in CDM.
@@
Our distinction. Previous studies have focused on question-side bias, while our work targets the Question Selection Module, mitigating bias from adaptive selection and directly addressing it at its source rather than applying post-hoc debiasing, thereby improving fairness and robustness in CAT.
$2.$ Bias Mitigation Methods
@@
While machine learning systems can relieve humans from tedious tasks and perform complex calculations at high speed, they are only as reliable as the data on which they are trained (barocas2016big). Since algorithms are not intentionally designed to incorporate bias, they often replicate or even amplify the spurious correlations present in real-world data, raising concerns regarding fairness and accountability (pedreshi2008discrimination; van2022overcoming). To address these issues, a variety of strategies have been proposed. Invariant Risk Minimization aims to learn environment-invariant predictors (arjovsky2019irm), while GroupDRO seeks to minimize worst-group risk but requires group labels (sagawa2019groupDRO). Reweighting strategies adjust sample importance to counter imbalance (japkowicz2000reweight), yet they only reshape optimization weights without enriching data diversity. In contrast, Mixup-based methods (zhang2018mixup) regularize models by interpolating samples and labels, with extensions such as Manifold Mixup (verma2019manifold) and Co-Mixup (kim2021co) improving hidden representation learning and pair selection. Building on this line of work, our method adopts a selective Mixup strategy to mitigate selection bias in CAT, leveraging its simplicity and effectiveness to enhance generalization.
@@ 3.1. Computerized Adaptive Testing
Computerized Adaptive Testing is composed of two main components: a Cognitive Diagnosis Module for proficiency estimation and a Question Selection Module. In the following, we provide a detailed description of the structure of these two components.
@@ 3.1.1. Cognitive Diagnosis Module
The Cognitive Diagnosis Module (CDM) estimates an examinee’s proficiency $\boldsymbol{\theta}$ by fitting their historical responses and predicting performance on unseen questions. Formally, the model outputs the probability $p\in[0,1]$ that an examinee answers a given item correctly. A representative neural-network-based CDM is NCDM (wang2020ncd), which leverages neural networks to capture complex interactions between examinees and items. The model incorporates several components: an examinee proficiency vector $\boldsymbol{\theta}\in(0,1)^{K}$, a question concept vector $\boldsymbol{Q}_{i}\in\{0,1\}^{1\times K}$ indicating the required knowledge concepts, a knowledge difficulty vector $\boldsymbol{h}^{diff}\in(0,1)^{K}$, a scalar discrimination parameter $h^{disc}\in(0,1)$, and trainable neural network parameters.
@@
The first interaction layer is inspired by the Multidimensional Item Response Theory (MIRT) (ackerman2014mirt) model, and is defined as:
@@
where $\circ$ denotes the element-wise product.
The intermediate layers consist of two fully connected layers followed by an output layer:
(2a) $\boldsymbol{f}_{1}$ $=\varphi(\boldsymbol{W}_{1}\boldsymbol{x}^{T}+\boldsymbol{b}_{1}),$
(2b) $\boldsymbol{f}_{2}$ $=\varphi(\boldsymbol{W}_{2}\boldsymbol{f}_{1}+\boldsymbol{b}_{2}),$
(2c) $p$ $=\varphi(\boldsymbol{W}_{3}\boldsymbol{f}_{2}+\boldsymbol{b}_{3}),$
where $\varphi$ denotes the activation function, typically chosen as the sigmoid. The monotonicity assumption requires that the probability of correctly answering an item increases monotonically with respect to each dimension of an examinee’s knowledge proficiency. To enforce this property, NCDM adopts a simple yet effective strategy by constraining all elements of $\boldsymbol{W}_{1}$, $\boldsymbol{W}_{2}$, and $\boldsymbol{W}_{3}$ to be positive.
@@
The cross-entropy loss is employed between the predicted probability $p\in[0,1]$ and the ground-truth response $y$:
(3) $\ell(y,p)=-\big(y\log p+(1-y)\log(1-p)\big),$
For simplicity, we denote the model output as $M(q;\boldsymbol{\theta},\zeta_{q},\zeta_{n})=p$. Here, $q$ denotes a question, $\boldsymbol{\theta}$ represents the examinee’s proficiency, $\zeta_{q}$ corresponds to the question parameters (e.g., concept vector, difficulty, discrimination), and $\zeta_{n}$ denotes the trainable neural network parameters. After training, the estimated proficiency vector $\boldsymbol{\theta}$ serves as the diagnostic result, characterizing the examinee’s proficiency across the $K$ knowledge concepts.
@@ 3.1.2. Question Selection Module
Existing data-driven Question Selection Modules typically employ a bi-level optimization framework (ghosh2021bobcat), where a selection network is trained on large-scale examinee data. The dataset $\mathcal{D}$ is partitioned into a support set $\mathcal{D}_{u}$ and a meta set $\mathcal{D}_{m}$, which are used to compute the loss or reward. As an illustrative example, we briefly introduce BOBCAT (ghosh2021bobcat). At each step $t$, the selection network $\pi(\cdot)$ chooses a question $q_{t}\in\mathcal{D}_{u}$ based on the examinee’s current state $\boldsymbol{s}_{t}$, which is encoded from their interaction history $r=\{(q_{1},y_{1}),\ldots,(q_{t-1},y_{t-1})\}$, where $y_{i}\in\{0,1\}$ indicates whether the response was correct or incorrect.
The state $\boldsymbol{s}_{t}$ is represented as a vector with length equal to the number of questions, where each entry takes values in $\{-1,0,1\}$, corresponding to incorrect, unattempted, and correct responses, respectively. Formally, the selection process is defined as:
(4) $q_{t}=\pi(\boldsymbol{s}_{t},\mathcal{D}_{u};\zeta_{s}),$
@@
To train the selection network, we adopt a bi-level optimization framework. The outer loop updates only the selection network $\zeta_{s}$, since the Question Selection Module is assumed to be the primary source of bias. The CDM is kept fixed, with parameters $(\zeta_{q},\zeta_{n})$ pre-trained and frozen, while the proficiency representation $\boldsymbol{\theta}$ is updated in the inner optimization. After training, the selection network generates the next question for each examinee based solely on their current response state. The overall bi-level optimization procedure for data-driven question selection can be summarized as follows, outlining the inner and outer optimization.
@@
Given a selected question sequence, the examinee’s proficiency $\boldsymbol{\theta}_{i}$ is estimated by minimizing the prediction loss on their responses in the support set $\mathcal{D}_{u}$, with fixed question parameters $(\zeta_{q},\zeta_{n})$. Through this process, $\boldsymbol{\theta}_{i}$ is updated to obtain the estimated proficiency vector $\boldsymbol{\theta}_{i}^{*}$:
@@
where $T$ denotes the number of selected questions.
@@
The selection network $\zeta_{s}$ is updated by minimizing the prediction loss on the meta set $\mathcal{D}_{m}$, using the fixed proficiency $\boldsymbol{\theta}_{i}^{*}$ obtained from the inner optimization and the pre-trained CDM parameters $(\zeta_{q},\zeta_{n})$:
@@
where NN denotes the number of examinees. In our framework, the key contribution lies in introducing Mixup-based data augmentation within this outer optimization stage, thereby enhancing the robustness of the selection network and mitigating bias.
Figure 3. An overview of our debiasing framework for Computerized Adaptive Testing. The method consists of two stages: (1) Cross-Attribute Examinee Retrieval, where midperformers from $\mathcal{D}_{m}^{\text{B}}$ are retrieved as neutral references for examinees in $\mathcal{D}_{m}^{\text{A}}$ and $\mathcal{D}_{m}^{\text{C}}$ based on proficiency similarity; and (2) Selective Mixup Regularization, where Mixup is applied between retrieved midperformers and their matched counterparts (with identical labels) to generate synthetic training samples. This augmentation reduces reliance on bias-aligned patterns, thereby mitigating bias in the Question Selection Module.